OpenStudy (anonymous):

helppppp What is the y-intercept in -4x - 6y = 13 ?

OpenStudy (anonymous):

y= mx+b b= y intercept

jimthompson5910 (jim_thompson5910):

or you can plug in x = 0 and solve for y -4x - 6y = 13 -4(0) - 6y = 13 ... ... ... y = ??

OpenStudy (anonymous):

19 ?

jimthompson5910 (jim_thompson5910):

-6y = 13 so y = ???

jimthompson5910 (jim_thompson5910):

-6y means -6 times y

jimthompson5910 (jim_thompson5910):

and you do NOT add 6 to both sides

OpenStudy (anonymous):

-78

jimthompson5910 (jim_thompson5910):

you don't multiply both sides by -6 either but you're getting closer

jimthompson5910 (jim_thompson5910):

-6 times y how do you undo the multiplication of -6?

OpenStudy (anonymous):

division

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so y = -13/6

OpenStudy (anonymous):

-2.1666666667

jimthompson5910 (jim_thompson5910):

you could use that or leave it as a fraction

OpenStudy (anonymous):

I put that as a decimal before and it was worng

jimthompson5910 (jim_thompson5910):

ok try entering it as a fraction then